OpenStudy (anonymous):

Evaluate the definite integral 4 ∫ 1/(x^2-6x+10) dx 2

OpenStudy (noelgreco):

Complete the square for the quadratic in the denominator.

OpenStudy (anonymous):

This is an ideal situation for using partial fractions. Have you used the partial fractions method for integrating before?

OpenStudy (anonymous):

@genius12, the denominator doesn't factor nicely though

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

denominator does not factor over integers, i would complete the square i think

OpenStudy (anonymous):

just noticed that lol. @SithsAndGiggles

OpenStudy (anonymous):

@shyllat, \[\begin{align*}x^2-6x+10&=x^2-6x+9+1\\ &=(x-3)^2+1 \end{align*}\] \[\int_2^4\frac{1}{(x-3)^2+1}~dx\] Familiar with trig substitution?

OpenStudy (anonymous):

ohhh ok yah i got it from here thank you!!